OpenStudy (anonymous):

during a rain storm the level of river rose 1 1/4 inches in 4 1/2 hours. How many inches did it rise in one hour? A.5 1/8 inches per hour B.5/36 inches per hour C.3 3/5 inches per hour D.3 5/8 inches per hour

OpenStudy (anonymous):

\[\frac{1\frac{1}{4}inch}{4\frac{1}{2}hours}=\frac{\frac{5}{4}inch}{\frac{9}{2}hour}=\frac{5}{4}inch \times \frac{2}{9}hours^{-1}=\frac{10}{36}=\frac{5}{18}inches/hour\]

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

thank you so much